26 February 2008

RHIC Running:  Attempts to reproduce the problem with the Blue abort kicker did not succeed.  There is currently no evidence for mistiming of the Blue abort kicker.  At this point, there is no good evidence that there was ever a problem with the abort kicker timing.

There was a short access yesterday to unblock a cooling water pump for the holding magnet in the JET.  PHENIX took the opportunity to swap out a LV Power Supply.  STAR/PP2PP also made an access to exchange a crate.

Polarization has been poor most of the night; <40%.  Some of this is due to reduced polarization out of the AGS which was down to 45% for at least one fill. 

Today, immedaitely after the 9:00AM meeting we will swap the orbits for the JET.  We will attempt to revert to the Run06 settings in AGS.  This may improve emittance blow-up, and improve polarization.  There will be work on polarization today. 

APEX will begin tonight at 8:00PM til 8:00AM Wednesday morning when the Maintenance Day will begin.  The primary task of APEX will be to revisit the beta* squeeze. 

Following the Maintenance Day on Thursday we will try to complete the nonlinear chromaticity correction. 

 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020